OpenStudy (ttp):

Simplify the expression 3 * (4 +(6-2) ^2)? 60 32 28 18

OpenStudy (anonymous):

\[3\times(4+(6-2)^2)?\]

OpenStudy (ttp):

yes:)

OpenStudy (shamil98):

3(4 + 16) =20(3) = 60

OpenStudy (anonymous):

A

OpenStudy (anonymous):

\[3\times(4+(4)^2) => 3 \times (4+16) => 3\times20 = 60\]